OpenStudy (anonymous):

Graph the line that passes through the points (-6, 8) and (5, -3).

OpenStudy (sydthekid913):

First you plot the points and find the slope Slope=rise over run

OpenStudy (sydthekid913):

The slope in this case is 1

OpenStudy (sydthekid913):

Then you find the y intercept

OpenStudy (sydthekid913):

Y-intercept = (0,x)

OpenStudy (sydthekid913):

The y intercept is 2

OpenStudy (sydthekid913):

You can probably figure out the rest for yourself

OpenStudy (sydthekid913):

:)

OpenStudy (triciaal):

@SydtheKid913 watch your signs the slope is change in y/ corresponding change in x slope =-3-8/5--6 = -11/11 = -1

OpenStudy (sydthekid913):

O whoops my bad :/ Thank you for that :)

OpenStudy (triciaal):

@sydthekid913 no problem @JohnDoemath40 can you continue ?

OpenStudy (anonymous):

ok my rise and run 11/11

OpenStudy (sydthekid913):

-

OpenStudy (triciaal):

@JohnDoemath40 you have to do the calculations in the same order hence corresponding -11/11 to get -1 as the slope

OpenStudy (triciaal):

now use the point-slope formula with one of the given points and the point (x, y)

OpenStudy (triciaal):

actually the question did not ask for the equation of the line to graph a line we need a minimum of 2 points we have the 2 points. plot the 2 points and connect them!

OpenStudy (triciaal):